• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:03:07 on June24,2002

    Entry number 86754

    keyword=Shift Summary

    Day Shift 6/24/02
    John LeRose: (SL)
    Gerfried Kumbartzki (TO)
    Pierre Moussiegt (AM)
    Hassan Ibrahim (PM)

    Running at kinematics Q2_j50 (All runs, unless otherwise indicated, are "production runs")
    Left Arm: P0 = 3.9682 19.312 deg
    Right Arm: P0 = 1.22760 66.269 deg
    Target: 15 cm LD2

    run # 1791 in progress at start of shift
    8:55 start run 1792
    9:56-10:05 Coda Problems run 1793 is junk. Reset ROCS and we're back in business
    10:06 start run 1794
    11:21 start run 1795 (after resetting ROC14)
    11:55 ended run 1795 starting ARC energy measurement
    12:55 ARC energy measurement done.
    E=4703.1 MeV +- 0.7 MeV (stat.) +- 0.9 MeV (Sys.) see: ELOG#186634
    13:08 MCC says 15 minutes before beam. They need to cycle the Moller quads.
    13:38 NOW MCC says they're ready to restore beam
    13:52 start run 1796 (Problem with ROC14!)
    14:10 start run 1798
    14:30 Q1 left tripped (He level fault)
    14:50 Q1 restored
    14:51 run 1799 started
    15:58 run 1800 started continues into next shift.
    Very good beam!